# Finite-sample Identification of Continuous-time Parameter-linear Systems

### v1

@@ 4.3 Bound on $\Delta\theta$
{pf} According to Lemma 9, $\operatorname{\mathbb{E}}\frac{\left\|v\right\|}{\sigma_{\text{min}}(R)}$ $\leq\frac{\epsilon_{N,m}^{1/2}}{\overline{\sigma}n^{\frac{1}{2m+3}}}$ According to Lemma 14, we can bound $\left\|u\right\|\operatorname{\mathbb{E}}\left\|S\right\|$ as the sum of $\left\|u\right\|s_{\text{head}}$ $=\frac{4\bar{U}\lambda}{\overline{\sigma}^{2}\sqrt{n^{\prime}}},$ $\left\|u\right\|s_{\text{body}}$ $=\frac{6C_{\phi}^{2\alpha}\bar{U}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}}+\frac{\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2}}\right],$ and $\left\|u\right\|s_{\text{tail}}$ $=\frac{4C_{\phi}^{2}\bar{U}\epsilon_{N}\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$ Using Lemmas 12 and 9 in Hölder’s inequality, $\operatorname{\mathbb{E}}\left\|S\right\|\left\|v\right\|$ $\leq\left(\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}\right)^{\alpha/2}\left(\operatorname{\mathbb{E}}\left\|v\right\|^{2/(2-\alpha)}\right)^{(2-\alpha)/2}$ $\leq\left(\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}\right)^{\alpha/2}$ $\quad\cdot\left(\frac{\epsilon_{N,m}^{1/2}(n^{\prime})^{1/2}}{n^{\frac{1}{2m+3}}}\right)$ $\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}$ $\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}.$
@@
which consolidates the two terms of Lemma 12, making the $\left\|S\right\|\left\|v\right\|$ estimate proportional to the $\left\|v\right\|/\sigma_{\text{min}}(R)$ estimate. We apply this choice to Thm. 15 to get the following.
Corollary 16
Under the choice $\lambda=\lambda^{*}$ of (51), $\operatorname{\mathbb{E}}\left\|\Delta\theta\right\|=O(n^{-\alpha/(2m+3)})$ as $n\to\infty$.
Appendix A Proof of Lemma 11
@@
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-2}$ $=\lambda^{2}+\sigma_{\text{min}}^{2}(A+D).$ (64)
@@
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-1}$ $\geq\frac{\lambda+\sigma_{\text{min}}(A+D)}{\sqrt{2}}.$ (67)
@@
Appendix B Proof of Lemma 13

### v3

@@ B.2 Bound on $\Delta\theta$
{pf}
According to Lemma 12,
$\operatorname{\mathbb{E}}\frac{\left\|v\right\|}{\sigma_{\text{min}}(R)}$ $\leq\frac{\epsilon_{N,m}^{1/2}}{\overline{\sigma}n^{\frac{1}{2m+3}}}$
According to Lemma 17, we can bound $\left\|u\right\|\operatorname{\mathbb{E}}\left\|S\right\|$ as the sum of
$\left\|u\right\|s_{\text{head}}$ $=\frac{4\bar{U}\lambda}{\overline{\sigma}^{2}\sqrt{n^{\prime}}},$
$\left\|u\right\|s_{\text{body}}$ $=\frac{6C_{\phi}^{2\alpha}\bar{U}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}}+\frac{\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2}}\right],$
and
$\left\|u\right\|s_{\text{tail}}$ $=\frac{4C_{\phi}^{2}\bar{U}\epsilon_{N}\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$
Using Lemmas 15 and 12 in Hölder’s inequality,
$\operatorname{\mathbb{E}}\left\|S\right\|\left\|v\right\|$ $\leq\left(\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}\right)^{\alpha/2}\left(\operatorname{\mathbb{E}}\left\|v\right\|^{2/(2-\alpha)}\right)^{(2-\alpha)/2}$
$\leq\left(\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}\right)^{\alpha/2}$
$\quad\cdot\left(\frac{\epsilon_{N,m}^{1/2}(n^{\prime})^{1/2}}{n^{\frac{1}{2m+3}}}\right)$
$\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}$
$\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}.$
@@
which consolidates the two terms of Lemma 15, making the $\left\|S\right\|\left\|v\right\|$ estimate proportional to the $\left\|v\right\|/\sigma_{\text{min}}(R)$ estimate. Applying the choice $\lambda=\lambda^{*}$ of (40) to Thm. 18 proves Theorem 8.
Appendix C Proof of Lemma 14
@@
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-2}$ $=\lambda^{2}+\sigma_{\text{min}}^{2}(A+D).$
@@
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-1}$ $\geq\frac{\lambda+\sigma_{\text{min}}(A+D)}{\sqrt{2}}.$
@@
Appendix D Proof of Lemma 16